(2t×siny+e to the power t×y to the power 3)dt+(t to the power 2×cosy+3×y to the power 2×e to the power t)dy=0
1
Expert's answer
2020-07-22T18:57:19-0400

D.E is

"(2t\sin(y)+e^ty^3)dt+(t^2\cos(y)+3y^2e^t)dy=0"

Now .divide both side of the above equation by "dt" ,thus we get

"(2t\sin(y)+e^ty^3)+(t^2\cos(y)+3y^2e^t)\frac{dy}{dt}=0\\
\implies 2t\sin(y)+t^2\cos(y)\frac{dy}{dt}+e^ty^3+3y^2e^t\frac{dy}{dt}=0\\"

Note that

"2t\sin(y)+t^2\cos(y)\frac{dy}{dt}=\frac{d}{dt}(t^2\sin(y))\\
e^ty^3+3y^2e^t\frac{dy}{dt}=\frac{d}{dt}(y^3e^t)"

Thus,

"2t\sin(y)+t^2\cos(y)\frac{dy}{dt}+e^ty^3+3y^2e^t\frac{dy}{dt}=0\\\implies 
\frac{d}{dt}(t^2\sin(y)+y^3e^t)=0\\
\implies \int d(t^2\sin(y)+y^3e^t)=\int0dt\\
\implies t^2\sin(y)+y^3e^t=c"